Prince Distribution Inc. has an unfunded postretirement benefit plan. Medical care and life insurance benefits are provided to employees who render 10 years service and attain age 55 while in service. At the end of 2018, Jim Lukawitz is 40. He was hired by Prince at age 35 (5 years ago) and is expected to retire at age 61. The expected postretirement benefit obligation for Lukawitz at the end of 2018 is 6$56,000 and $59,000 at the end of 2019 Calculate the accumulated postretirement benefit obligation at the end of 2018 and 2019 and the service cost for 2018 and 2019 as pertaining to Lukawitz. (Enter your answers in whole dollar amount.) APBO Service Cost 2018 + 2019 < Prey 11 of 14 … Next > 26
Answer #1

Attribution Period = (55-35) = 20 Year

Accumulated postretirement benefit obligation at the end of 2018 = EPBO*5/20

Accumulated postretirement benefit obligation at the end of 2018 = 56000*5/20

Accumulated postretirement benefit obligation at the end of 2018 = $ 14000

Accumulated postretirement benefit obligation at the end of 2019 = EPBO*6/20

Accumulated postretirement benefit obligation at the end of 2019 = $ 59,000*6/20

Accumulated postretirement benefit obligation at the end of 2019 = $ 17,700

Service cost for 2018 = EPBO*1/20

Service cost for 2018 = 56000*1/20

Service cost for 2018 = 2,800

Service cost for 2019 = EPBO*1/20

Service cost for 2019 = 59,000*1/20

Service cost for 2019 = 2,950